# Knowledge-based Consistency Testing of Large Language Models

### v1

@@ Abstract
In this work, we systematically expose and measure the inconsistency and knowledge gaps of Large Language Models (LLMs). Specifically, we propose an automated testing framework (called KonTest) which leverages a knowledge graph to construct test cases. KonTest probes and measures the inconsistencies in the LLM’s knowledge of the world via a combination of semantically-equivalent queries and test oracles (metamorphic or ontological oracle). KonTest further mitigates knowledge gaps via a weighted LLM model ensemble. Using four state-of-the-art LLMs (Falcon, Gemini, GPT3.5, and Llama2), we show that KonTest generates 19.2% error inducing inputs (1917 errors from 9983 test inputs). It also reveals a 16.5% knowledge gap across all tested LLMs. KonTest’s mitigation method reduces LLM knowledge gap by 32.48%. Our ablation study further shows that GPT3.5 is not suitable for knowledge-based consistency testing because it is only 60%-68% effective in knowledge construction.
@@
Sai Sathiesh Rajan SUTD, Singapore sai_rajan@mymail.sutd.edu.sg                        Ezekiel Soremekun Royal Holloway, University of London ezekiel.soremekun@rhul.ac.uk
Sudipta Chattopadhyay SUTD, Singapore sudipta_chattopadhyay@sutd.edu.sg
@@ 1 Introduction
KonTest leverages a knowledge graph for consistency testing (see Figure 2). It first automatically extracts a set of entities and entity relationships from the knowledge graph. This is then used to systematically generate (semantically-equivalent) queries and create test cases for the subject LLM under test (SUT) across various settings (e.g., atomic query vs. sequential query). The test case generation involves minimal, one-time effort for each type of entity relation considered (e.g., only two templates for 6730 test cases) and such templates can be reused for testing arbitrary LLMs. Additionally, KonTest automatically mitigates knowledge gaps by leveraging the likelihood of LLM inconsistencies to construct a weighted model ensemble. To the best of our knowledge, KonTest is the first systematic approach for automatically generating consistency tests for assessing LLMs.
@@
    KonTest automatically mitigates knowledge gaps via a weighted model ensemble via the likelihood of LLM inconsistencies. The mitigation technique of KonTest reduces knowledge gaps by 32.48% (section 5).
@@ 2 Overview
Figure 3: Example queries and templates in KonTest
@@ 3.1 Knowledge Graph based Test Generation
Test Generation: Given a knowledge path $\mathbb{KG}$, KonTest leverages Algorithm 1 to exhaustively generate queries relating to all possible pairs of entities present in the path. To this end, KonTest first finds the relation, $\mathbb{KG}_{rel}$, between a pair of nodes. KonTest then generates a query, ${Query}$, using $\mathbb{KG}_{rel}$ with the aid of the template shown in Figure 3 (9 in Algorithm 1). KonTest also generates the mutated query, ${Query_{M}}$, with another template (12 in Algorithm 1). These templates are dependent on the relation between the two nodes in question. For instance, Figure 3 shows the set of queries and mutated queries (i.e., $Query_{list}$ and $MutQuery_{list}$) generated from one particular path and along with the respective templates used to generate them. We note that developers can implement additional templates easily if they are interested in querying different relations. Moreover, creation of these templates is a one-time process and the cost incurred is minimal since the number of different types of relationships is small in comparison to the number of entities.
@@ 3.3 Knowledge Gap Mitigation
KonTest then uses the relative weights to calculate the final ensemble score by first multiplying the relative weights with the results for each response and summing them. In particular, it considers a positive and negative responses to be one and zero respectively. If the ensemble score is above the midpoint (0.5), the ensemble response for the relation is considered to be positive. To determine the effectiveness of our weighting, we compare it to simple majority voting – an ensemble that does not account for the likelihood of inconsistencies for each SUT (see RQ3).
@@ 4 Evaluation Setup
  • RQ1 Effectiveness: How effective is KonTest in exposing consistency errors in LLMs?
@@
  • RQ3 Knowledge Gap Mitigation: How effective is KonTest’s weighted ensemble technique in improving knowledge coverage? How does it compare to a simple majority voting? Does it generalize to a new template?
@@
Knowledge Graph: We rely on Wikidata’s knowledge graph Vrandečić and Krötzsch (2014) and Wikidata’s SPARQL query service to extract information pertinent to the two tested domains: places and music. We choose these domains as inconsistent knowledge about places may have adversarial consequences in navigation use cases of LLMs Mapbox . Concurrently, inconsistencies in music domain may inaccurately capture the ownership of copyrighted materials (e.g., music albums). For the places domain, we take the list of countries ranked by (nominal) GDP per capita and select the top ten countries as our initial entity list IMF . We then iteratively build our knowledge graph by finding “administrative divisions" that are “located in the administrative territorial entity" of the parent entity. For the music domain, we take the top 200 musical acts since 2000 as our initial entity list Chart (2000). We then find the set of “albums" that have the parent entity as a “performer". We also find the set of songs and singles that are “part of" the album being considered. Once the graph is fully constructed, we randomly select 50 leaf nodes for each domain and extract the full path associated with the selected leaf nodes.
@@
Knowledge Gap Mitigation: We evaluate the effectiveness of KonTest’s ensemble method by determining relative weights assigned to each SUT. We restrict the set of relations considered for this process such that a relation that yields an invalid answer in the query generation phase for any of the SUTs is not used in subsequent computations. We then perform five-fold cross validation on the remaining set of relations. Concretely, we hold out 20% of the relations for evaluation and use the consistency errors associated with the rest of the relations in the computation of the weights. We introduce a third template to evaluate whether KonTest’s mitigation generalizes to unseen templates.
@@
Table 3: Effectiveness of KonTest using a knowledge graph. (Meta.: Metamorphic, Onto.: Ontological)
LLM (Subject)
Valid Test
Executions (%)
Errors (%)
Meta. Onto. All Meta. Onto. All
Places Falcon 1322 (89.9) 267 (91.4) 1589 (90.2) 226 (17.1) 25 (9.4) 251 (15.8)
Gemini 1470 (100) 292 (100) 1762 (100) 161 (11.0) 40 (13.7) 201 (11.4)
GPT3.5 1468 (99.9) 292 (100) 1760 (99.9) 400 (27.2) 17 (5.8) 417 (23.7)
Llama2 1351 (91.9) 268 (91.8) 1619 (91.9) 355 (26.3) 27 (10.1) 382 (23.6)
Total 5611 (95.4) 1119 (95.8) 6730 (95.5) 1142 (20.4) 109 (9.7) 1251 (18.6)
Music Falcon 715 (97.9) 94 (97.9) 809 (97.9) 131 (18.3) 3 (3.2) 134 (16.6)
Gemini 730 (100) 96 (100) 826 (100) 124 (17.0) 7 (7.3) 131 (15.9)
GPT3.5 730 (100) 96 (100) 826 (100) 181 (24.8) 5 (5.2) 186 (22.5)
Llama2 696 (95.3) 96 (100) 792 (95.9) 206 (29.6) 9 (9.4) 215 (27.1)
Total 2871 (98.3) 382 (99.5) 3253 (98.5) 642 (22.4) 24 (6.3) 666 (20.5)
Total 8482 (96.4) 1501 (96.7) 9983 (96.4) 1784 (21.0) 133 (8.9) 1917 (19.2)
@@ 5 Evaluation Results
Domain Relations Knowledge Gap (%)
Template 1 Template 2 Template 3 All (%Impr.)
RQ1 Effectiveness: We found that KonTest is effective in exposing consistency errors in LLMs: 19.2% of valid test executions result in consistency errors. Table 3 also shows that metamorphic errors (21.0%) are more prevalent than ontological errors (8.9%). We attribute the lower error rate of ontological errors/inputs to the high complexity of our ontological oracle (see Table 1). Table 3 demonstrates that metamorphic errors are common across all LLMs. In particular, the tested LLMs are highly inconsistent when asked the same question in a different manner (see Table 1). In addition, we observe that a large model size does not necessarily lead to a smaller error rate than a smaller model, e.g., GPT3.5 exhibits a metamorphic error rate of 27.2% for the places entities as compared to a metamorphic error rate of 17.1% for Falcon.
@@
RQ3 Knowledge Gap Mitigation: Results show that KonTest’s mitigation technique reduces the SUT’s knowledge gap by up to 39.30%. Table 4 shows that KonTest reduces the knowledge gap for all templates by 32.48%. We found that the simple majority voting ensemble worsens the SUTs’ knowledge gap by up to 23.74%. We also observed that the mitigation performance of KonTest generalizes to an unseen template (template three (3)). While the mitigation performance of KonTest is slightly better on the new template than the original templates the simple majority performs worse on an unseen template. These results demonstrate the generalizability of KonTest’s mitigation and the efficacy of our weighted ensemble. We attribute the performance of KonTest’s mitigation to the likelihood and distribution of inconsistencies: Firstly, KonTest’s weighted ensemble is 42.72% more effective than the simple majority voting. Secondly, the distribution of inconsistencies shows that only 1.1% (15 out of 1330) of errors are found in all four SUTs while 65.7% (874 out of 1330) of errors are unique to a single SUT.
The ensemble method of KonTest effectively reduces the SUT’s knowledge gap by 32.48%.
@@
Knowledge Construction: Table 6 demonstrates that GPT3.5 is not a reliable knowledge constructor. GPT3.5 is only able to identify 68% of the relations present in the original knowledge graph. We believe this is because LLMs are generally intended to be a conversational engine, rather than a knowledge database Pan et al. (2023). These results show the importance of knowledge graphs in KonTest and suggests that LLMs are not a reliable replacement for knowledge graphs.
@@
Error Detection: Table 8 highlights that while 52.6% of the metamorphic errors detected by GPT3.5 were detected correctly, approximately 40.1% of the errors were false positives  [^2] Note that some errors detected by GPT3.5 cannot be automatically validated by our oracles (UNK in Table 8), as it excludes responses that do not begin with yes/no. . For ontological error detection, we provided GPT3.5 all possible queries and corresponding responses for each knowledge path. GPT3.5 was then asked to detect any inconsistency in these responses. Unlike KonTest, GPT3.5 is unable to detect multiple inconsistencies in a knowledge path. Hence, for a fair comparison in this experiment, we count ontological errors for KonTest at the granularity of a knowledge path (i.e., at most one error per knowledge path). Since validating GPT3.5 responses is not straightforward in this case, we manually validated the responses. We observed that the number of real ontological errors detected by GPT3.5 is comparable to the ontological errors detected by KonTest, but it also had a high false positive rate with nearly 63.7% of errors being misclassified.
@@ 7 Conclusion
In this paper, we propose KonTest where the key intuition is to distill (a subset of) facts from a knowledge graph, which, are subsequently used to generate queries and formulate test cases for detecting a variety of consistency errors. Our evaluation reveals realistic consistency errors across state-of-the-art LLMs. Moreover, KonTest opens opportunities to investigate LLMs through the lens of their knowledge gaps, which, in turn is indicated as part of our framework. This helps designers and end users to understand and mitigate the effect of such knowledge gaps e.g., via prompt engineering or fine tuning. In future, we aim to investigate automated mitigation of consistency errors by leveraging the KonTest framework. We provide our code and experimental data in the following:
http://anonymous.4open.science/r/Kontest
@@ 8 Limitations and Threats to Validity
Internal Validity: This refers to the threat that our implementation of KonTest performs its intended knowledge graph-based test generation. We conduct several manual and automated tests, as well as inspection of sampled outcomes of KonTest to ensure the correctness of our approach. Our ablation study (RQ4) further allows to probe the correctness of the sub-step of KonTest versus using GPT3.5. We experimentally verify that over 90% of the entities we evaluate against existed in Wikidata prior to 2020. We also find that under 15% of the errors found by KonTest are linked to these entities. In addition, we also note that the SUT ought to answer the question in a consistent manner regardless of whether the entity existed prior to the corresponding knowledge cutoff date.
External Validity: The main threat to external validity of this work is the generalizability of KonTest and findings to LLMs, knowledge graphs, templates and entities beyond the ones used in this work. For instance, we mitigate the LLM generalizability threat by employing state-of-the-art off-the-shelf, mature, open model weights LLMs (LLaMA2 and FALCON), as well as commercial LLMs (such as GPT3.5 and Gemini). Notably, our entity selection and template construction may be limited to our experimental setting. However, we demonstrate the applicability of KonTest by using two different domains with multiple relations and entity types. We note that we employ few-shot prompting to query the LLMs, thus, our findings may not generalise to other prompting techniques. Finally, KonTest employs Wikidata, a well-maintained knowledge graph that is popularly used in both academia and industry Peters et al. (2019).
@@
Knowledge Graph Completeness and Soundness: We note that the knowledge graph is an incomplete snapshot of the real-world. In our evaluation, KonTest only tests for facts (positive tests) derived from the knowledge graph. While it is also possible to use KonTest to test for incorrect relation (negative test), the validation of such test results is challenging due to the incompleteness of knowledge graph. Moreover, we only tests 50 paths from this graph. However, these concerns do not affect our findings, as we are certain about the errors found within the tested subset of the knowledge graph. Finally, knowledge of the world naturally evolves over time and the knowledge graphs do not evolve at the same pace Pan et al. (2023). To mitigate this, we use a widely used knowledge graph Vrandečić and Krötzsch (2014).

### v2

@@ Abstract
In this work, we systematically expose and measure the inconsistency and knowledge gaps of Large Language Models (LLMs). Specifically, we propose an automated testing framework (called KonTest) which leverages a knowledge graph to construct test cases. KonTest probes and measures the inconsistencies in the LLM’s knowledge of the world via a combination of semantically-equivalent queries and test oracles (metamorphic or ontological oracle). KonTest further mitigates knowledge gaps via a weighted LLM model ensemble. Using four state-of-the-art LLMs (Falcon, Gemini, GPT3.5, and Llama2), we show that KonTest generates 19.2% error inducing inputs (1917 errors from 9979 test inputs). It also reveals a 16.5% knowledge gap across all tested LLMs. A mitigation method informed by KonTest’s test suite reduces LLM knowledge gap by 32.48%. Our ablation study further shows that GPT3.5 is not suitable for knowledge-based consistency testing because it is only 60%-68% effective in knowledge construction.
@@ 1 Introduction
KonTest leverages a knowledge graph for consistency testing (see Figure 2). It first automatically extracts a set of entities and entity relationships from the knowledge graph. This is then used to systematically generate (semantically-equivalent) yes/no queries and create test cases for the subject LLM under test (SUT) across various settings (e.g., atomic query vs. sequential query). The test case generation involves minimal, one-time effort for each type of entity relation considered (e.g., only two templates for 6730 test cases) and such templates can be reused for testing arbitrary LLMs. KonTest’s test suite can additionally be used to mitigate knowledge gaps by leveraging the likelihood of LLM inconsistencies to construct a weighted model ensemble. To the best of our knowledge, KonTest is the first systematic approach for automatically generating consistency tests for assessing LLMs.
@@
    We propose a technique that mitigates knowledge gaps via a weighted model ensemble via the likelihood of LLM inconsistencies obtained from KonTest’s test suite. The mitigation technique reduces knowledge gaps by 32.48% (section 5).
@@ 2 Overview
Figure 3: Example queries and templates for the "Places" domain in KonTest
@@ 3.1 Knowledge Graph based Test Generation
Test Generation: Given a knowledge path $\mathbb{KG}$, KonTest leverages Algorithm 1 to exhaustively generate queries relating to all possible pairs of entities present in the path. To this end, KonTest first finds the relation, $\mathbb{KG}_{rel}$, between a pair of nodes. KonTest then generates a query, ${Query}$, using $\mathbb{KG}_{rel}$ with the aid of the template shown in Figure 3 (9 in Algorithm 1). KonTest also generates the mutated query, ${Query_{M}}$, with another template (12 in Algorithm 1). These templates are dependent on the relation between the two nodes in question. For instance, Figure 3 shows the set of queries and mutated queries (i.e., $Query_{list}$ and $MutQuery_{list}$) generated from one particular path and along with the respective templates used to generate them.
We note that developers can implement additional templates easily if they are interested in querying different relations. Moreover, creation of these templates is a one-time process and the cost incurred is minimal (<10 LoC) since the number of different types of relationships is small in comparison to the number of entities. We further note that our templates are directly applicable to other domains with similar relations. For instance, our templates for the "Places" domain can be reused for the "Food" domain since both domains have a "contains" type relation.
@@ 3.3 Knowledge Gap Mitigation
We then compute the final ensemble score by first multiplying the relative weights with the results for each response and summing them. In particular, we consider positive and negative responses to be one and zero respectively. If the ensemble score is above the midpoint (0.5), the ensemble response for the relation is considered to be positive. To determine the effectiveness of our weighting, we compare it to simple majority voting – an ensemble that does not account for the likelihood of inconsistencies for each SUT (see RQ3).
@@ 4 Evaluation Setup
  • RQ1 Effectiveness: How effective is KonTest in exposing consistency errors in LLMs? Are KonTest results stable?
@@
  • RQ3 Knowledge Gap Mitigation: How effective is the proposed weighted ensemble technique in improving knowledge coverage? How does it compare to a simple majority voting? Does it generalize to a new template?
@@
Knowledge Graph: We rely on Wikidata’s knowledge graph Vrandečić and Krötzsch (2014) and Wikidata’s SPARQL query service to extract information pertinent to the two tested domains: places and music. We choose these domains as inconsistent knowledge about places may have adversarial consequences in navigation use cases of LLMs Mapbox . Concurrently, inconsistencies in the music domain may inaccurately capture the ownership of copyrighted materials (e.g., music albums). For the places domain, we take the list of countries ranked by (nominal) GDP per capita and select the top ten countries as our initial entity list IMF . We then iteratively build our knowledge graph by finding “administrative divisions" that are “located in the administrative territorial entity" of the parent entity. For the music domain, we take the top 200 musical acts since 2000 as our initial entity list Chart (2000). We then find the set of “albums" that have the parent entity as a “performer". We also find the set of songs and singles that are “part of" the album being considered. Once the graph is fully constructed, we randomly select 50 leaf nodes for each domain and extract the full path associated with the selected leaf nodes.
@@
Knowledge Gap Mitigation: We evaluate the effectiveness of an ensemble method informed by KonTest’s test suite by determining the relative weights assigned to each SUT. We restrict the set of relations considered for this process such that a relation that yields an invalid answer in the query generation phase for any of the SUTs is not used in subsequent computations. We then perform five-fold cross validation on the remaining set of relations. Concretely, we hold out 20% of the relations for evaluation and use the consistency errors associated with the rest of the relations in the computation of the weights. We introduce a third template to evaluate whether our mitigation scheme generalizes to unseen templates.
@@ 5 Evaluation Results
Table 3: Effectiveness of KonTest using a knowledge graph.
LLM (Subject) Valid Test Executions (%) Errors (%)
Metamorphic Ontological All Metamorphic Ontological All
atomic
sequential
-intra
sequential
-inter
All Types atomic
sequential
-intra
sequential
-inter
All Types
Places Falcon 254 (86.4) 530 (90.1) 538 (91.5) 1322 (89.9) 267 (91.4) 1589 (90.2) 50 (19.7) 24 (4.5) 152 (28.3) 226 (17.1) 25 (9.4) 251 (15.8)
Gemini 294 (100) 588 (100) 588 (100) 1470 (100) 292 (100) 1762 (100) 51 (17.3) 6 (1.0) 104 (17.7) 161 (11.0) 40 (13.7) 201 (11.4)
GPT3.5 293 (99.7) 588 (100) 587 (99.8) 1468 (99.9) 292 (100) 1760 (99.9) 61 (20.8) 201 (34.2) 138 (23.5) 400 (27.2) 17 (5.8) 417 (23.7)
Llama2 266 (90.5) 543 (92.3) 542 (92.2) 1351 (91.9) 268 (91.8) 1619 (91.9) 94 (35.3) 52 (9.6) 209 (38.6) 355 (26.3) 27 (10.1) 382 (23.6)
Total 1107 (94.1) 2249 (95.6) 2255 (95.9) 5611 (95.4) 1119 (95.8) 6730 (95.5) 256 (23.1) 283 (12.6) 603 (26.7) 1142 (20.4) 109 (9.7) 1251 (18.6)
Music Falcon 141 (96.6) 287 (98.3) 287 (98.3) 715 (97.9) 94 (97.9) 809 (97.9) 24 (17.0) 38 (13.2) 69 (24.0) 131 (18.3) 3 (3.2) 134 (16.6)
Gemini 146 (100) 292 (100) 292 (100) 730 (100) 96 (100) 826 (100) 39 (26.7) 8 (2.7) 77 (26.4) 124 (17.0) 7 (7.3) 131 (15.9)
GPT3.5 146 (100) 292 (100) 292 (100) 730 (100) 96 (100) 826 (100) 40 (27.4) 85 (29.1) 56 (19.2) 181 (24.8) 5 (5.2) 186 (22.5)
Llama2 137 (93.4) 283 (96.9) 276 (94.5) 696 (95.3) 92 (95.8) 788 (95.4) 46 (33.6) 63 (22.3) 97 (35.1) 206 (29.6) 9 (9.8) 215 (27.3)
Total 570 (97.6) 1154 (98.8) 1147 (98.2) 2871 (98.3) 378 (98.4) 3249 (98.3) 149 (26.1) 194 (16.8) 299 (26.1) 642 (22.4) 24 (6.3) 666 (20.5)
Total 1677 (95.3) 3403 (96.7) 3402 (96.6) 8482 (96.4) 1497 (96.5) 9979 (96.4) 405 (24.2) 477 (14.0) 902 (26.5) 1784 (21.0) 133 (8.9) 1917 (19.2)
RQ1 Effectiveness: We found that KonTest is effective in exposing consistency errors in LLMs: 19.2% of valid test executions result in consistency errors. Table 3 also shows that metamorphic errors (21.0% = 1784/8482) are more prevalent than ontological errors (8.9% = 133/1497). We attribute the lower error rate of ontological errors/inputs to the high complexity of our ontological oracle (see Table 1). Table 3 demonstrates that metamorphic errors are common across all LLMs. In particular, the tested LLMs are highly inconsistent when asked the same question in a different manner (see Table 1). In addition, we observe that a large model size does not necessarily lead to a smaller error rate than a smaller model, e.g., GPT3.5 exhibits a metamorphic error rate of 27.2% for the places entities as compared to a metamorphic error rate of 17.1% for Falcon.
In addition, we validate the stability of our results by repeating our experiments four additional times. We find that the open-source models with frozen weights (Falcon and Llama2) yielded identical results when compared to the initial experiments. However, the closed-source models exhibited (up to 6%) lower error rates on the subsequent iterations with Gemini exhibiting a less than one percent change in error rate. We attribute this to changes in the underlying models OpenAI as the additional experiments were performed approximately eight (8) months after the initial experiments. Furthermore, we find that the error rates for all four subsequent iterations are fairly similar indicating that the results are consistent when repeated within a short duration of time (within a few days). In particular, we find that Gemini results do not vary at all. Lastly, we also examine the stability of our results by executing both atomic and sequential queries separately. For instance, consider Q1 and Q6a (Table 1) and the following additional oracle: LLM(Q1)=LLM(Q6a). We found that only GPT3.5 exhibited inconsistent results for this oracle and further noted that the absolute number of errors found was negligible (<1%). These experiments demonstrate the stability of our results.
@@
Domain Relations Knowledge Gap (%)
Template 1 Template 2 Template 3 All (%Impr.)
RQ3 Knowledge Gap Mitigation: Results show that our proposed mitigation technique reduces the SUT’s knowledge gap by up to 39.30%. Table 5 shows that our technique reduces the knowledge gap for all templates by 32.48%. We found that the simple majority voting ensemble worsens the SUTs’ knowledge gap by up to 23.74%. We also observed that the mitigation performance of our technique generalizes to an unseen template (template three (3)). While the performance of our mitigation technique is slightly better on the new template than the original templates the simple majority performs worse on an unseen template. On the whole, we find that our weighted ensemble is 42.72% ((323-185)/323) more effective than simple majority voting at reducing the knowledge gap. These results demonstrate the generalizability of our proposed mitigation technique and the efficacy of our weighted ensemble. We attribute the performance of our technique to the likelihood and distribution of inconsistencies discovered by KonTest (Figure 4). Figure 4 shows that only 1.1% (15 out of 1330) of errors are found in all four SUTs while 65.7% (874 out of 1330) of errors are unique to a single SUT. This suggests that an ensembling scheme, informed by the relative performance of the SUTs, reduces knowledge gap.
The proposed ensemble method effectively reduces the SUT’s knowledge gap by 32.48%.
Figure 4: Venn Diagram of the consistency errors for each SUT.
@@
Knowledge Construction: Table 6 demonstrates that GPT3.5 is not a reliable knowledge constructor. GPT3.5 is only able to identify 68% of the relations present in the original knowledge graph. We believe this is because LLMs are generally intended to be a conversational engine, rather than a knowledge database Pan et al. (2023). These results show the importance of knowledge graphs in KonTest and suggest that LLMs are not a reliable replacement for knowledge graphs.
@@
Error Detection: Table 8 highlights that while 52.6% of the metamorphic errors detected by GPT3.5 were detected correctly, approximately 40.1% of the errors were false positives  [^2] Note that some errors detected by GPT3.5 cannot be automatically validated by our oracles (UNK in Table 8), as it excludes responses that do not begin with yes/no. . For ontological error detection, we provided GPT3.5 all possible queries and corresponding responses for each knowledge path. GPT3.5 was then asked to detect any inconsistency in these responses. Unlike KonTest, GPT3.5 is unable to detect multiple inconsistencies in a knowledge path. Hence, for a fair comparison, we count ontological errors for KonTest at the granularity of a knowledge path (i.e., at most one error per knowledge path). Since validating GPT3.5 responses is not straightforward in this case, we manually validated the responses. We observed that the number of real ontological errors detected by GPT3.5 is comparable to the ontological errors detected by KonTest, but it also had a high false positive rate with nearly 63.7% of errors being misclassified.
@@ 7 Conclusion
In this paper, we propose KonTest where the key intuition is to distill (a subset of) facts from a knowledge graph, which are subsequently used to generate queries and formulate test cases for detecting a variety of consistency errors. Our evaluation reveals realistic consistency errors across state-of-the-art LLMs. Moreover, KonTest opens opportunities to investigate LLMs through the lens of their knowledge gaps, which, in turn, is indicated as part of our framework. This helps designers and end users to understand and mitigate the effect of such knowledge gaps e.g., via prompt engineering or fine tuning. In future, we aim to investigate automated mitigation of consistency errors by leveraging the KonTest framework. We provide our code and experimental data in the following:
https://github.com/sparkssss/KonTest
@@ 8 Limitations and Threats to Validity
Internal Validity: This refers to the threat that our implementation of KonTest performs its intended knowledge graph-based test generation. We conduct several manual and automated tests, as well as inspection of sampled outcomes of KonTest to ensure the correctness of our approach. Our ablation study (RQ4) further allows us to probe the correctness of the sub-step of KonTest versus using GPT3.5. We experimentally verify that over 90% of the entities we evaluate against existed in Wikidata prior to 2020. We also find that under 15% of the errors found by KonTest are linked to these entities. In addition, we also note that the SUT ought to answer the question in a consistent manner regardless of whether the entity existed prior to the corresponding knowledge cutoff date.
External Validity: The main threat to external validity of this work is the generalizability of KonTest and findings to LLMs, knowledge graphs, templates and entities beyond the ones used in this work. We mitigate the LLM generalizability threat by employing state-of-the-art off-the-shelf, mature, open model weights LLMs (LLaMA2 and FALCON), as well as commercial LLMs (such as GPT3.5 and Gemini). Similarly, our entity selection and template construction may be limited to our experimental setting. However, we demonstrate the applicability of KonTest by using two different domains with multiple relations and entity types. However, we note that KonTest might not be easily adapted to information that cannot be encapsulated within a knowledge graph. In addition, we employ few-shot prompting to conduct our ablation study, and our findings might not generalize to other prompting techniques. Finally, KonTest employs Wikidata, a well-maintained knowledge graph that is popularly used in both academia and industry Peters et al. (2019).
@@
Knowledge Graph Completeness and Soundness: We note that the knowledge graph is an incomplete snapshot of the real-world. In our evaluation, KonTest only tests for facts (positive tests) derived from the knowledge graph. While it is also possible to use KonTest to test for incorrect relation (negative test), the validation of such test results is challenging due to the incompleteness of the knowledge graph. Moreover, we only test 50 paths from this graph. However, these concerns do not affect our findings, as we are certain about the errors found within the tested subset of the knowledge graph. Finally, knowledge of the world naturally evolves over time and the knowledge graphs do not evolve at the same pace Pan et al. (2023). To mitigate this, we use a widely used knowledge graph Vrandečić and Krötzsch (2014).